two insulating small spheres are rubbed against each other and are keptat a distance of 4m apart. if they attract each other with a force of 3.5 N then calculate charge on each sphere.

Grade:12th pass

1 Answers

Arun
6 years ago
Spheres have the same in magnitude and opposite signed charges q1 = − q2 = q;
Force evaluated due to Coulomb law:F = k * q2/r2;
⇒ q = r * (F/k)1/2 = 8 * 10-5C.  (r = 4 m, F = 3.5 N)
where k = 9 * 109 N * m2 / C2
Number of electrons equal to the ration between charge of sphere and single electron charge:
Ne = q/e = 5 * 1014 , where e = 1.6 * 10-19 C
Answer: q = 8 * 10-5 C , Ne = 5 * 1014 .